MOA-II Obsevation Log
Return to index
Previous day
Next day
DATE: 2016/02/06
OBSERVER: Koshimoto Paul
CONDITION: Clear
COMMENT: chip 1 is noisy in the 300 sec dark (D13477) but it won't be fixed for 2hours and the noise level is low
Compressor gas pressure: 1.8 MPa
Expected Observation Time: START 22:26 END 05:39
Go to the Seeing Log
Go to the Observation Log
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 52427-52435 | 9 | 54 | 3240 |
LMC | 12555-12573 | 19 | 55 | 16500 |
SMC | - | 0 | 0 | 0 |
TRANSIT | - | 0 | 0 | 0 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 28 | 109 | 19740 (05h29m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
5 | 3 | 7 | 7 | 5 | 0 | 1 | 2 | 4 | 4 | 2 | 1 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
2 | 4 | 0 | 0 | 1 | 1 | 1 | 1 | 1 | 0 | 2 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 19 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 18 |
Field13 | Field14 | Field15 | Field16 |
0 | 18 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
TRANSIT FIELD STATISTICS
SPECIAL TARGET NAMEBR>
CHECK SHEET
Time | Status | Weather |
22:26 | waiting | Clear |
23:00 | observing | Clear |
00:00 | observing | Clear |
01:00 | observing | Clear |
02:00 | observing | Clear |
03:00 | observing | Clear |
04:00 | observing | Clear |
05:00 | observing | Clear |
05:39 | observing | Clear |
CCD TEMPERATURE
Previous day
Next day

VACUUM
Previous day
Next day

TELESCOPE TEMPERATURE
Previous day
Next day

Return to index
Previous day
Next day